Sol Summary – February 25th

Crew 275 Sol Summary Report 25Feb2023

Sol: 13

Summary Title: Field Day for Crew 275

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

After almost two weeks at the MDRS, all the crewmembers are getting more and more efficient in all their tasks. This gives a little bit more time to relax between activities, which is appreciated by everyone. Today was the first geology EVA and it was a success. It started by a first stop at the atmospheric instruments site to change the batteries and retrieve the data, before heading to Kissing Camel Ridge W. The EVA Crew had the opportunity to test an instrument to easily measure both the magnetic susceptibility and the conductivity in real field conditions.

On the campus, everyone is now very happy to be able to follow the past and current temperature, humidity, and luminosity in all the buildings. Human factors experiments have continued, while our crew journalist has started working on our mid-rotation video. We are all looking forward to seeing the results!

As we are reaching the end of this second week, I also decided to spend some time to see each crewmember individually and talk about how they felt so far. I know that the third quarter of any long-duration challenge (like a Marathon, for instance) is the most difficult. It is also the case for analog missions like ours, and I want to make sure that everyone is in the best condition for this period.

Look Ahead Plan:

Today we have changed the batteries of the atmospheric instruments. This means that tomorrow can be a real rest day without any EVAs. Next week, we would like to plan 3 additional EVAs dedicated to photogrammetry: One for the mapping of the area, one for the exploration with a 2D map, and the third one for the exploration with a 3D render. This time the target area would be Candor Chasma.

Anomalies in work: None.

Weather: Sunny, low -4°C/25°F, high 6°C/43°F.

Crew Physical Status: Optimal.

EVA: EVA #12 was successful. All the objectives have been accomplished and the collected samples will soon be analyzed.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site